### v1

@@ III-D Interactive Feedback
1:$prompt$ = $\text{template}(\Delta g_{t},\delta g_{t},\Delta s_{t},s_{t})$triangleright Predefined Template
2:$result$ = call GPT-4 API with $prompt$
@@
Interactive perception. It should be noted that occlusion and estimation errors may arise during the perception of the first observation. To address these challenges, we propose a model that leverages interactive observations to enhance operation (see Algo. 1). Consider two distinct observations, $I_{rgbd}^{0}$ and $I_{rgbd}^{t}$, captured during an interaction. A part’s motion can be detected between these two observations. We subsequently employ the mask of the moved part to deduce the actual joint state during the interaction. Applying RANSAC [9] for outlier removal and Umeyama algorithm [50], we estimate the rotation and translation between the two frames of the part’s point cloud $X_{\text{part}}^{0},X_{\text{part}}^{t}$. This computation subsequently allows us to ascertain the joint state $j_{t}$ and the current part movement states $s_{t}$.
Tracking and planning update. Throughout the interaction, tracking of both the target gripper and part states is maintained. If significant discrepancies arise, the planner has the discretion to select from one of four states (see Algo. 2): ”continue”, ”transition to the next step”, ”halt and replan”, or ”success”. For instance, if the gripper is set to rotate 60 degrees along a joint, yet the door only opens to 15 degrees, the LLM planner would opt to ”halt and re-plan”. This interaction tracking model ensures that the LLM remains cognizant of developments during interactions, permitting strategy adjustments and, where necessary, recovery from unforeseen setbacks.
@@
Mean Opinion Score
GAPartNet
VLM
Ours
VLM
Ours
Part description $\uparrow$ 3.8 4.1 6.3 7.2 9.6
Part accuracy $\uparrow$ 8.7 4.0 6.5 6.9 9.5
Part state precision $\uparrow$ 0.0 3.7 5.6 7.7 7.8
Object&scene description $\uparrow$ 0.0 2.7 6.2 6.4 8.0
Interaction-related info.$\uparrow$ 0.3 7.3 7.6 7.6 8.7
Overall performance. $\uparrow$ 2.6 4.4 6.4 7.2 8.7
TABLE I: User study results for scene description. We solicited volunteers to assess the quality of scene descriptions. Participants were tasked with evaluating the following aspects: (1) Performance in the overall part description. (2) Accuracy in the enumeration and naming of parts. (3) Precision in part-state depiction. (4) Depiction of objects and the overall scene. (5) Description of interaction-related information. (6) Overall performance.
Figure 6: Demonstrations of real-world results. Top: ”turn on the blender” and ”turn on/off the machine”. Our method accurately understands the part semantics and actions that are not aligned. Bottom: ”open the microwave”. The mechanical structures of the microwave prevent the robot from directly pulling open the door but instead require the button to be pressed, resulting in a failure. However, our interactive feedback model can detect the failure and recognize that it should try pressing the button instead and subsequently complete the task.
Part Perception In-distribution Unseen states Unseen objects Unseen categories Accuracy AP@50 mAP AP@50 mAP AP@50 mAP AP@50 mAP PointGroup[28] 69.70 60.58 69.54 60.48 58.26 46.29 24.57 19.40 SoftGroup[52] 69.59 60.54 70.02 59.59 59.20 47.10 28.18 22.50 AutoGPart[35] 66.81 57.63 67.60 56.69 55.30 43.50 26.24 20.38 GAPartNet[14] 81.42 72.55 80.80 71.73 63.18 53.94 36.39 27.40 PartGroundedSAM [30, 40] 73.73 61.75 72.36 62.02 66.25 38.64 41.59 28.45 Ours 83.04 72.23 82.39 71.91 72.17 58.04 47.69 34.57 TABLE II: Part perception results. We have curated a novel evaluation dataset for part perception, enriched with more comprehensive data. Our method is benchmarked against 3D point cloud-based techniques, including PointGroup[28], SoftGroup[52], AutoGPart[35], and GAPartNet[14]. In alignment with the 2D branch of our approach, we also employ SAM[30] and DINOv2[40] to establish a 2D-centric baseline. For evaluation, we adopt AP@50 and mAP as our primary metrics.
Figure 7: Visual plots of the part perception results. Our method consistently achieved the highest AP@50 score for each level of generalization.
Obj. Cat. Microwave StorageFurniture Cabinet KitchenPot Remote Blender
@@
$*{}^{*}$[24](%) - - 13.0 - - 15.0 - - 14.2 - - -
$*{}^{*}$[14](%) 87.5 75.5 58.6 53.3 76.9 81.3 50.0 66.3 73.8 - - -
Ours(%) 98.0 96.0 94.1 83.3 80.0 95.0 79.3 89.6 83.1 85.7 60.7 42.9
TABLE III: Language-guided manipulation results. Comparison between our method with VoxPoser [24] and GAPartNet[14]. We evaluate 12 tasks with 6 different articulated objects. ”-” means the task is not implemented by the baseline. $*{}^{*}$ shares the same execution policy as our method, while $*{}^{*}$ is adapted from the unofficial version of the authors. Obj.Cat.=Object Category
Obj. Cat. # Tasks # Initial state # Target state # Instruction
@@
TABLE IV: Benchmark statistics. Detailed statistics for our language-guided manipulation benchmark. For each object category, we created tasks that were randomly initialized. Obj.Cat.=Object Category
IV Experiments
IV-A Scene description.
@@ IV Experiments
We evaluate the scene descriptions generated by five methods: (1) GAPart Model: detection results are translated into natural language descriptions through a human-crafted format; (2)(3) VLM (Bard, GPT-4V), craft descriptions from the input image using Bard and GPT-4V; (4)(5) Our methods (based on Bard and GPT-4V). In the user study, participants are provided an image and the descriptions generated by the three methods at each time, and they are asked to rate the descriptions using a scale of 0-10 (with 10 being the best) according to the 6 metrics. 17 participants have been invited, and each person is asked to rate about 40 scenes on average. Table I shows the results of this user study. We found that domain-specific models can hardly provide useful object and scene-level descriptions due to their lack of general knowledge. On the other hand, VLM provides rich context but lacks accuracy in domain-related information. Our method benefits from both of them and achieves much better performances by joining their strengths.
IV-B Part perception.
@@
We also introduce a new benchmark for part perception evaluation. In comparison with GAPartNet, our benchmark is more comprehensive and suitable for manipulation tasks. For instance, we incorporate a greater number of objects with closed parts, accounting for 12.5%, which is not included in GAPartNet[14]. To enable the evaluation of generalizations at different levels, the test data are divided into 4 subsets: in-distribution, unseen (articulation) states, unseen objects, and unseen object categories. We use the average precision AP@50 and mAP as evaluation metrics for part segmentation, which are widely adopted in prior 3D semantic/instance segmentation benchmarks such as ScanNet[6] and GAPartNet[14],
@@
Table II presents the results of our method in comparison to various baselines for part perception. We consider GAPartNet[14], a modified version of PointGroup[28], SoftGroup[52], and AutoGPart [35] as our primary baselines. Additionally, we adapted GroundedSAM, rebranding it as PartGroundedSAM for this context. The performance metrics in Table II clearly indicate that our approach surpasses other baselines. Notably, we observed that methods based on 3D tend to under-perform on out-of-domain data, whereas 2D-centric methods yield subpar results for parts. Our methodology derives advantages from both the 2D and 3D realms, resulting in optimal performance and superior generalization capabilities.
IV-C Manipulation in Simulation Environments
We conducted our simulations using the SAPIEN environment [56] and designed 12 language-guided articulated object manipulation tasks. A comprehensive breakdown of these tasks, along with detailed statistics, is presented in Table III IV. For each category of Microwave, StorageFurniture and Cabinate, we devised 3 tasks including open from the initial open and close state and closing from the initial open state. The remaining tasks for KitchenPot, Remote and Blender are Open the lid, Press the button and Turn on the blender respectively. For each task, we carefully curated more than 5 distinct objects sourced from the GAPartNet dataset [14] that were well-suited for the respective actions. For example, we selected 5 different models of Microwaves from GAPartNet, chosen specifically for tasks like pulling the door open and pushing the door closed. Likewise, we chose 20 different StorageFurniture objects and 10 Blenders, each tailored to their corresponding tasks. To assess the robustness of our part interaction module, we conducted over 20 trials for each task. During each trial, we introduced randomization in both camera position and initial joint states, ensuring the variability of scenarios. Specifically, for the task Close the door, the initial positions of the articulated object doors were randomized within the range of $(30,60)$ degrees. To facilitate fair comparisons, we leveraged pre-trained weights from GAPartNet to identify actionable parts within the scene. Subsequently, we applied the same motion generation policy as our pipeline. The results of our experiments are summarized in Table III, showcasing the superior performance of our methods across nearly all tasks. The key drivers of our success can be attributed to our enhanced perception capabilities, which benefit from the fusion of the specialized model from GAPartNet and the generalist model of LLM.
IV-D Real-World Manipulation
In our real-world experiments, we establish an experimental setup with the UFACTORY xArm 6 and several different articulated objects for manipulation. Fig. 6 shows some featured results, and more results can be found in the supplementary video and our website. Fig. 6 top left shows a challenging case of activating a blender. The top part of the blender is perceived as a container for containing juices but functions as a button to be pressed down. Our framework effectively links its semantic and action understandings and successfully executes the task. Fig. 6 top right is an emergency stop button for robots, which requires a press (down) to halt an operation and a rotation (up) to restart it. With the auxiliary input of the user manual, our method completes both of these two tasks. Fig. 6 bottom shows a rather challenging case where the microwave door cannot be directly pulled open. Instead, it requires first pressing the button to initiate a slight door opening. In this case, our method first tries the most straightforward solution of pulling the door and fails. The interactive feedback module then detects this failure and replans a second strategy that adapts to the current environment, which finally completes the task.
V Conclusions
@@
In this paper, we introduce a novel framework for language-guided manipulation of articulated objects. Bridging the understanding of object semantics and actionability at the part level, we are able to ground language-implied actions to executable manipulations. Throughout our framework, we also study the combination of general-purpose large vision/language models and domain-specialist models for enhancing the richness and correctness of network predictions. better handle these tasks and achieve state-of-the-art performance. We demonstrate our strong generalizability across diverse object categories and tasks. We also provide a new benchmark for language-instructed articulated-object manipulations.

### v2

@@ IV-D Interactive Feedback
1:$prompt$ = $\text{template}(\Delta g_{t},\delta g_{t},\Delta s_{t},s_{t})$
2:$result$ = call VLM with $prompt$
@@
Algorithm 2 Interactive perception
      Output joint axis $j_{t}$ and movement $s_{t}$
2:$j_{t},s_{t}=\text{GeometryInference}(R_{t},t_{t})$
@@
Interactive perception. It should be noted that occlusion and estimation errors may arise during the perception of the first observation. To address these challenges, we propose a model that leverages interactive observations to enhance operation (see Algo. 2). Consider two distinct observations, $I_{rgbd}^{0}$ and $I_{rgbd}^{t}$, captured during an interaction. A part’s motion can be detected between these two observations. We subsequently employ the mask of the moved part to deduce the actual joint state during the interaction. Applying RANSAC [9] for outlier removal and Umeyama algorithm [48], we estimate the rotation and translation between the two frames of the part’s point cloud $X_{\text{part}}^{0},X_{\text{part}}^{t}$. This computation subsequently allows us to ascertain the joint state $j_{t}$ and the current part movement states $s_{t}$. Finally, we send the joint state and part movement state back to the global planner (see Sec. IV-B, thereby completing the main loop.
V Experiments
@@
Category Microwave StorageFurniture Cabinet KitchenPot Remote Blender
@@
VoxPoser[24] - - 13.0 - - 15.0 - - 14.2 - - -
GAPartNet[13] 87.5 75.5 58.6 53.3 76.9 81.3 50.0 66.3 73.8 - - -
Ours 98.0 96.0 94.1 83.3 80.0 95.0 79.3 89.6 83.1 85.7 60.7 42.9
TABLE I: Success rates (%) under language instructions. Comparison between our method with VoxPoser [24] and GAPartNet[13]. We evaluate 12 tasks with 6 different articulated objects. ”-” means the task is not implemented by the baseline. GAPartNet shares the same execution policy as our method, while VoxPoser is adapted from the unofficial version of the authors.
Category #Task #Init. state #Tgt. state #Instruct.
@@
TABLE II: Benchmark statistics. For each object category, we created tasks that were randomly initialized.
Perception Scene Description <1s
Part Detection 13s
Decision Instruction Interpretation <10s
Global Planninng <1s
Execution Part Grounding <5s
Motion Planning 45s
Feedback Interactive Perception <1s
TABLE III: Average runtime of each module.
V-A Simulation Experiments
Setup
We conducted our simulations using the SAPIEN environment [55] and designed 12 language-guided articulated object manipulation tasks. An example task is given in Fig. 6. A comprehensive breakdown of these tasks, along with detailed statistics, is presented in Table I II. For each category of Microwave, StorageFurniture and Cabinate, we devised 3 tasks including opening from the initial open and close state and closing from the initial open state. The remaining tasks for KitchenPot, Remote and Blender are Open the lid, Press the button and Turn on the blender respectively. For each task, we carefully curated more than 5 distinct objects sourced from the GAPartNet dataset [13] that were well-suited for the respective actions. For example, we selected 5 different models of Microwaves from GAPartNet, chosen specifically for tasks like pulling the door open and pushing the door closed. Likewise, we chose 20 different StorageFurniture objects and 10 Blenders, each tailored to their corresponding tasks. To assess the robustness of our part interaction module, we conducted over 20 trials for each task. During each trial, we introduced randomization in both camera position and initial joint states, ensuring the variability of scenarios. Specifically, for the task Close the door, the initial positions of the articulated object doors were randomized within the range of $(30,60)$ degrees. To facilitate fair comparisons, we leveraged pre-trained weights from GAPartNet to identify actionable parts within the scene. Subsequently, we applied the same motion generation policy as our pipeline.
Results
The results of our experiments are summarized in Table I, showcasing the superior performance of our methods across nearly all tasks. The key drivers of our success can be attributed to our enhanced perception capabilities, which benefit from the fusion of the specialized model from GAPartNet and the generalist model of LLM. The average runtime breakdown of each module in the SAGE system is shown in Table. III.
Figure 7: Explanations of real-world results. Top: ”turn on the blender” and ”turn on/off the machine”. Our method accurately understands the part semantics and actions that are not aligned. Bottom: ”Open the microwave”. The mechanical structures of the microwave prevent the robot from directly pulling open the door but instead require the button to be pressed, resulting in a failure. However, our interactive feedback model can detect the failure and recognize that it should try pressing the button instead and subsequently complete the task.
Figure 8: Real-robot results. We show the keyframes of various tasks in our experiments. More results can be found in the supplementary materials.
V-B Real-Robot Experiments
Setup
In our real-world experiments, we establish an experimental setup with the UFACTORY xArm 6 and several different articulated objects for manipulation.
Results
Fig. 7 shows our results on real-robot execution, and more results can be found in the supplementary materials. Three challenging cases are highlighted in Fig. 7 with detailed explanations and intermediate outputs. The top left is a blender whose top part is perceived as a container for containing juices but functions as a button to be pressed down. Our framework effectively links its semantic and action understandings and successfully executes the task. The top right shows an emergency stop button for robots, which requires a press (down) to halt an operation and a rotation (up) to restart it. With the auxiliary input of the user manual, our method completes both of these two tasks. The bottom shows a rather challenging case where the microwave door cannot be directly pulled open. Instead, it requires first pressing the button to initiate a slight door opening. In this case, our method first tries the most straightforward solution of pulling the door and fails. The interactive feedback module then detects this failure and informs the global planner to replan a second strategy that adapts to the current environment, finally completing the task.
V-C Generalizable Visual Perception
MOS $\uparrow$ GAPartNet
VLM
Ours
VLM
Ours
Part description 3.8 4.1 6.3 7.2 9.6
Part accuracy 8.7 4.0 6.5 6.9 9.5
Part state precision 0.0 3.7 5.6 7.7 7.8
Object & scene descr. 0.0 2.7 6.2 6.4 8.0
Interaction info. 0.3 7.3 7.6 7.6 8.7
Overall performance 2.6 4.4 6.4 7.2 8.7
TABLE IV: User study results for scene description. We solicited volunteers to assess the quality of scene descriptions. Participants were tasked with evaluating the following aspects: (1) Performance in the overall part description. (2) Accuracy in the enumeration and naming of parts. (3) Precision in part-state depiction. (4) Depiction of objects and the overall scene. (5) Description of interaction-related information. (6) Overall performance.
Part Perception In-distribution Unseen states Unseen objects Unseen categories
Accuracy AP@50 mAP AP@50 mAP AP@50 mAP AP@50 mAP
PointGroup[27] 69.70 60.58 69.54 60.48 58.26 46.29 24.57 19.40
SoftGroup[50] 69.59 60.54 70.02 59.59 59.20 47.10 28.18 22.50
AutoGPart[33] 66.81 57.63 67.60 56.69 55.30 43.50 26.24 20.38
GAPartNet[13] 81.42 72.55 80.80 71.73 63.18 53.94 36.39 27.40
PartGroundedSAM [29, 38] 73.73 61.75 72.36 62.02 66.25 38.64 41.59 28.45
Ours 83.04 72.23 82.39 71.91 72.17 58.04 47.69 34.57
TABLE V: Part perception results. We have curated a novel evaluation dataset for part perception, enriched with more comprehensive data. Our method is benchmarked against 3D point cloud-based techniques, including PointGroup[27], SoftGroup[50], AutoGPart[33], and GAPartNet[13]. In alignment with the 2D branch of our approach, we also employ SAM[29] and DINOv2[38] to establish a 2D-centric baseline. For evaluation, we adopt AP@50 and mAP as our primary metrics.
Figure 9: Part perception results. Our method consistently achieved the highest AP@50 score for each level of generalization.
As described in Sec. IV, we design knowledge fusion mechanisms to join the force VLMs and small domain-specific models both in our context comprehension and part perception. In this section, we evaluate our intermediate results on scene description (Sec. IV-A) and part perception (Sec. IV-C) and show that our generalizable perception modules achieve a better balance between generalization and exactness compared to existing methods or its ablated versions.
Scene description (Sec. IV-A)
@@ V-C Generalizable Visual Perception
We evaluate the scene descriptions generated by five methods: (1) GAPart Model: detection results are translated into natural language descriptions through a human-crafted format; (2)(3) VLM (Bard, GPT-4V), craft descriptions from the input image using Bard and GPT-4V; (4)(5) Our methods (based on Bard and GPT-4V). In the user study, participants are provided an image and the descriptions generated by the three methods at each time, and they are asked to rate the descriptions using a scale of 0-10 (with 10 being the best) according to the 6 metrics. 17 participants have been invited, and each person is asked to rate about 40 scenes on average.
Table IV shows the results of this user study. We found that domain-specific models can hardly provide useful object and scene-level descriptions due to their lack of general knowledge. On the other hand, VLM provides rich context but lacks accuracy in domain-related information. Our method benefits from both of them and achieves much better performances by joining their strengths.
Part perception (Sec. IV-C)
@@
In our part perception task, we utilize a single RGBD image as input to predict part-related information, including part semantic segmentation, pose and state estimation. To evaluate the methods, we introduce a new benchmark for part perception tasks. In comparison with GAPartNet, our benchmark is more comprehensive and suitable for manipulation tasks. For instance, we incorporate a greater number of objects with closed parts, accounting for 12.5%, which is not included in GAPartNet[13]. To enable the evaluation of generalizations at different levels, the test data are divided into 4 subsets: in-distribution, unseen (articulation) states, unseen objects, and unseen object categories. We use the average precision AP@50 and mAP as evaluation metrics for part segmentation, which are widely adopted in prior 3D semantic/instance segmentation benchmarks such as ScanNet[6] and GAPartNet[13],
@@
Table V presents the results of our method in comparison to various baselines for part perception. We consider GAPartNet[13], a modified version of PointGroup[27], SoftGroup[50], and AutoGPart [33] as our primary baselines. Additionally, we adapted GroundedSAM, rebranding it as PartGroundedSAM for this context. The performance metrics in Table V indicate that our approach surpasses other baselines. Notably, we observed that methods based on 3D tend to under-perform on out-of-domain data, whereas 2D-centric methods yield subpar results for parts. Our methodology derives advantages from both the 2D and 3D realms, resulting in optimal performance and superior generalization capabilities.
VI Conclusions
@@
In this paper, we introduce a novel framework for language-guided manipulation of articulated objects. Bridging the understanding of object semantics and actionability at the part level, we can ground language-implied actions to executable manipulations. Throughout our framework, we also study the combination of general-purpose large vision/language models and domain-specialist models for enhancing the richness and correctness of network predictions. better handle these tasks and achieve state-of-the-art performance. We demonstrate our strong generalizability across diverse object categories and tasks. We also provide a new benchmark for language-instructed articulated-object manipulations.
Limitations and future work
Since the SAGE system is built upon existing large vision-language models, it suffers from the drawbacks of demanding massive data and massive computing power. Although in our case we use APIs provided by GPT-4V, the long inference time and high cost could be a concern. On the other hand, we rely on the expert model GAPartNet to detect parts and its prior knowledge to manipulate the parts. The policy is not always optimal by blindly following the pre-defined end-effector trajectories. Such a motion-planning-based execution policy is not as responsive as a reinforcement learning or imitation learning agent. One future exploration direction could be fine-tuning existing large models to directly output the desired low-level action of the end-effector to increase responsiveness.